OpenStudy (hiyathere_20):

Evaluate A. 64 B. 5 C. 4 D. 262,144

Awolflover1 (awolflover1):

Evaluate what?

Awolflover1 (awolflover1):

\[8^\frac{ 2 }{ 3 }? \]

OpenStudy (hiyathere_20):

yes

Awolflover1 (awolflover1):

hmm what did you get?

OpenStudy (hiyathere_20):

idk how to solved this. I need help . u tell me ?(:

Awolflover1 (awolflover1):

\[\sqrt{8^2}\]

OpenStudy (hiyathere_20):

64?

Awolflover1 (awolflover1):

no no

Awolflover1 (awolflover1):

\[2^2\]

OpenStudy (hiyathere_20):

4

Awolflover1 (awolflover1):

there we go

OpenStudy (hiyathere_20):

okayyyyy???? so how did I get 64?

Awolflover1 (awolflover1):

\[8^2 =8\times8\]